OpenStudy (nuccioreggie):

The number line shows the distance in meters of two skydivers, A and B, from a third skydiver located at point X: A horizontal number line extends from negative 3 to positive 3. The point labeled as A is at negative 2.5, the point 0 is labeled as X, and the point labeled B is at 0.5. Write an expression using subtraction to find the distance between skydivers A and B. (5 points) Show your work and solve for the distance using additive inverses. (5 points)

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

your problem can be modeled by this subsequent drawing: 

OpenStudy (nuccioreggie):

Ok

OpenStudy (michele_laino):

so, the request distance d, is given by the subsequent formula: d=B-A=0.5-(-2.5) So, what is: 0.5-(-2.5)=...?

OpenStudy (nuccioreggie):

3

OpenStudy (nuccioreggie):

Now i think we have to Show your work and solve for the distance using additive inverses.

OpenStudy (michele_laino):

that's right! Your answer is right!

OpenStudy (nuccioreggie):

Ok

OpenStudy (michele_laino):

we can write this: \[d = B - A = \frac{1}{2} - \left( { - \frac{5}{2}} \right) = \frac{1}{2} + \frac{5}{2} = \frac{{1 + 5}}{2} = ...?\] please continue

OpenStudy (nuccioreggie):

Ok can youwrite it the other way so i can be able to write it in the answer box please

OpenStudy (michele_laino):

I think that you have to write: 3

OpenStudy (nuccioreggie):

No i have to Show your work and solve for the distance using additive inverses.

OpenStudy (michele_laino):

we can write this: first way: d=0.5-(-2.5)= =(3-2.5)-(-2.5)= =3-2.5+2.5 second way: d=0.2-(-2.5)= =0.5-(0.5-3)= 0.5-0.5+3

OpenStudy (nuccioreggie):

For the first one i got. 3 - 2 3

OpenStudy (nuccioreggie):

That the first way am i correct

OpenStudy (michele_laino):

I think that both ways are correct!

OpenStudy (nuccioreggie):

O so do i need to answer them or no??

OpenStudy (michele_laino):

I think that you have to choose only one of them

OpenStudy (nuccioreggie):

Ok but do i have to answer one of them? Do you want me to a nswer one of them

OpenStudy (nuccioreggie):

??

OpenStudy (michele_laino):

I think that you can choose the first one!

OpenStudy (michele_laino):

namely the first way!